40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu May 14 17:41:10 2015, rana, Update, PEM, weather station and Guralp maintenance GurPost_150514.pngsecretWeatherTrends.png
    Reply  Sat May 16 21:40:29 2015, rana, Update, PEM, Guralp maintenance GurPost_150516.png
       Reply  Thu May 28 10:42:19 2015, Steve, Update, PEM, bad Guralp X-cable Gurs180dXbad.pngswappedGurs.png
          Reply  Thu May 28 15:10:44 2015, rana, Update, PEM, Seismic Confusion 
    Reply  Tue May 26 11:05:10 2015, Steve, Update, PEM, worked around ETMY seism. ETMYseismic.pngcut.jpg
Message ID: 11291     Entry time: Thu May 14 17:41:10 2015     Reply to this: 11295   11324
Author: rana 
Type: Update 
Category: PEM 
Subject: weather station and Guralp maintenance 

Today Steve and I tried to recenter the Guralps. The breakout box technique didn't work for us, so we just turned the leveling screws until we got the mass position outputs within +/-50 mV for all DoF as read out by the breakout box.

Some points:

  1. GUR1 is at the ETMY (E/W arm) and GUR2 is at the X-end (South arm)
  2. The SS containers are good and make a good seal.
  3. We had to replace the screws on the granite slab interface plate. The heads were too big to allow the connector to snap into place.
  4. The Guralps had been left way, way off level and the brass locking screws were all the way up. We locked them down after leveling today. Steve was blaming Cathy(?).
  5. The GUR1_Z channel now looks good - see the summary pages for the before and after behavior. My mistake; the low frequency is still as bad as before.
  6. GUR2 X/Y still look like there is no whitening or if the masses are stuck or the interface box is broken.
  7. When we first powered them up, a few of the channels of both seismometers showed 100-200 Hz oscillations. This has settled down after several minutes.

 

The attachment shows the 6 channels after our work. You can see that GUR2_X/Y still look deadish. I tried wiggling the cables at the interface box and powering on/off, but no luck. Next, we swap cables.

Tried to bring the weather station back to life, but no luck. The unit on the wall is alive and so is the EPICS IOC (c1pem1). But there is apparently no communication between them. telnet into c1pem and the error message repeating at the prompt is:

Weather Monitor Output: NO COMM

Might be related to the flaky connector situation that Liz and I found there a couple summers ago, but I tried jiggling and reseating that one with no luck. Looks like it stopped working around 8 PM on March 24, 2014. That's the same time as a ~30s power outage, so perhaps we just need some more power cycling? Tried hitting the reset button on the VME card for c1pem1, but didn't change anything.

Let's try power cycling that crate (which has c1pem1, c0daqawg, and some GPS receiver)...nope - no luck.

Also tried power cycling the weather box which is near the BS chamber on the wall. This didn't change the error message at the c1pem1 telnet prompt.

Attachment 1: GurPost_150514.png  13 kB  | Hide | Hide all
GurPost_150514.png
Attachment 2: secretWeatherTrends.png  17 kB  Uploaded Thu May 14 19:13:57 2015  | Hide | Hide all
secretWeatherTrends.png
ELOG V3.1.3-